OpenStudy (anonymous):

DOES ANYONE KNOW HOW TO SOLVW THESE>??? #1) cos x(sec x – cos x) #2) (1 + cot x)^2

OpenStudy (anonymous):

i think you were asked to use distributive property here.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for number 1, in distributing, note that, secx and cosx are reciprocal of are other.

OpenStudy (anonymous):

so, cosx*(secx) = ____?

OpenStudy (campbell_st):

you have \[\cos(x)(\frac{1}{\cos(x)} - \cos(x)) = 1 - \cos^2(x) = \sin^2(x)\]

OpenStudy (campbell_st):

this is a binomial expansion of a perfect square \[(1 + \cot(x))^2 = 1 + 2\cot(x) + \cot^2(x)\] use \[\cos^2(x) + 1 = \csc^2(x)\] so the solution is \[\csc^2(x) + 2\cot(x)\]